For the gas N2 calculate the volume in mL if 189 g were under a pressure of 543 torr at 40.0 oC.

A sample of gas contains 62.0 moles at 327 K in 4.00 L of volume. What is the pressure (in atm) of the gas?

What is the density (in mg/mL) of a H2S gas at 12.9 atm and 137 K

Place the following gases in order of increasing average molecular speed at 25 oC:

Slowest → Fastest energy

CO, Br2, HF
Answer #1

Ans. w = 189 g , M = 28g/mol , P = 543 torr , T = 40 celsius = 313 K

It is given that

n = 62 moles , T = 327 K, V = 4.00L

PV = n R T

P = n R T / V = 62 * 0.082 * 327 / 4 = 415.61 atm

density = P M / R T = 12.9 * 34 / 0.082 * 137 = 39.04 mg/mL

av. speed is inversely proporational to the square root of molar mass

hence order is:- Br2 < CO < HF